# Higher-order shortest paths in hypergraphs

### v2

@@ INTRODUCTION
Networks, collections of nodes and their interactions, are one of the primary tools used to study complex systems, allowing us to describe collective, emergent system properties which can not be captured by looking at the individual units in isolation newman2018networks . A typical case is the emergence of global connectivity from local interactions, with many real-world networks characterised by the emergence of large connected components molloy1995critical . Moreover, such systems display low diameter and are hence dubbed ‘small-worlds’ wattsstrogatz1998smallworld , supporting a relational structure able to support both efficient system-wide communications latora2001efficient ; estrada2008communicability and make them easy to navigate kleinberg2000navigation . The efficient structural connectivity of these systems has a profound impact on their functionality, with the presence of shortcuts increasing the speed of contagion and the emergence of cooperation santos2005epidemic , and the ability of a system to synchronise barahona2002synchronization . For all these reasons, efficiently computing shortest paths in graphs is a problem that has attracted enormous interest in the research community for many decades dijkstra2022note , and it continues to do so.
@@
Networks have classically modelled interactions through links, describing relations between pairs of entities only, even though in many real-world systems interactions simultaneously involve multiple nodes battiston2020honetworks . Recently, a wide variety of structural descriptors have been extended to account for the presence of such higher-order interactions, including algorithms for motif discovery lotito2022homotifs and analysis mann2021randomgraphs , community detection eriksson2022flow ; ruggeri2023community , centrality measures benson2019three , as well as network filtering musciotto2021detecting and reconstruction young2021hypergraph procedures. Non-dyadic interactions generate new dynamical behaviours and collective phenomena battiston2021physics ; bick2023higher , from contagion iacopini2019simplicial ; dearruda2020social to synchronisation skardal2020higher ; millan2020explosive ; zhang2023higher and evolutionary games alvarez2021evolutionary ; civilini2024explosive . More recently, temporality has also been considered in higher-order networks, from burstiness cencetti2021temporal and temporal-topological correlations ceria2023temporal , to Markovian petri2018simplicial ; digaet2024percolation and non-Markovian gallo2024higher ; iacopini2024temporal structural models.
Despite these advances, characterising shortest paths and connectivity in systems with higher-order interactions remains an open problem. Recently, efforts have been devoted to characterise the concepts of distance vasilyeva2023distancesho and walks aksoy2020hypernetwork and networks in networks with non-dyadic ties, as well as proposing efficient algorithms gao2014hoshortestpath to extract shortest paths in hypergraphs, limiting the analysis to static systems.
In this work, we provide the first systematic investigation of the effect of higher-order interactions on system connectivity across a variety of real-world datasets. We quantify the contribution of non-dyadic ties for shortest path length, computing length distribution, average path size and the fraction of purely dyadic segments in each path, and show that all such features are compatible with a simple null-model which preserves the higher-order degree distribution. Next, we consider time-varying interactions, and extend the concept of higher-order path and components to temporal hypergraphs. Our analysis of multiple real-world systems reveals the crucial role of higher-order interactions to ensure efficient connectivity in temporal higher-order networks, characterises differences between topologically shortest and temporally fastest paths, and shows how the observed non-trivial empirical patterns can not be reproduced with simple randomised null-models. Our work provides new insights into the connectivity properties and the structural organisation of real-world hypergraphs, both for static and time-varying systems.
@@ RESULTS
To study shortest paths in temporal hypergraphs, we collected 25 publicly available datasets of real-world temporal systems with higher-order interactions. These hypergraphs describe social interactions from a broad range of domains such as households and schools, hospitals, workspaces and conferences, as well as political interactions between individuals in Congress and even contact data from baboons
Dataset $V$ $E_{HO}$ $E_{DY}$ $|e_{\max}^{HO}|$ $T$ $dt$
@@
In simple graphs, relational information is described by edges, which encode pairwise interactions between pairs of nodes only. A path is an ordered, non-repeating sequence of nodes, where each subsequent step occurs along a particular edge. Similarly, a hypergraph is a collection of nodes and hyperedges which encode interactions among an arbitrary number of nodes. The size of a hyperedge is defined as the number of nodes contained therein. A hyperpath or a higher-order path in a hypergraph is defined as an ordered, non-repeating sequence of nodes, where each subsequent step occurs between nodes within the same hyperedge.
The length of a higher-order path between any two nodes $i$ and $j$, $\ell_{ij}$, is defined as the number of nodes it traverses minus 1. Intuitively, this may be thought of as the number of ‘steps’ taken. If no such path exists between 2 nodes, we set $\ell_{ij}=\infty$. Two nodes are connected if a path of finite length exists between them. While a pair of nodes can be connected by multiple distinct paths, the shortest path between any two nodes in a static network is defined as a path of minimum length, which might not necessarily be unique.
Figure 1: The shortest hyperpath between nodes B and K has length $\ell=3$ and average size $\langle{s}\rangle=(4+5+2)/3$.
@@
Clearly, hyperpaths in higher-order networks will traverse hyperedges of varying sizes, where each segment of a hyperpath traverses a hyperedge of a potentially different size. As such, one may attach to each individual segment of a hyperpath its associated size $s$, defined as the size of the hyperedge that is traversed during that particular ‘step’. In some cases, such a step may be contained in multiple hyperedges, in which case we choose the hyperedge of smallest size to define its size. Paths in higher-order networks contain rich information and one may calculate the average size $\langle{s}\rangle$ of a path as the average size across traversed segments. For example, the average size of a path allows one to determine how much of a path is composed of genuinely higher-order interactions. The minimum size of a path is $\langle{s}\rangle=2$, which occurs when the hyperpath consists of purely pairwise interactions.
@@
We begin by considering static hypergraphs where we neglect information about the temporal nature of the interactions (see Methods for details). As an illustrative example, in Figure 2, we show results for the social hypergraph from the well-known Copenhagen network study, which describes social interactions over 4 weeks between 576 university students on a university campus.
@@
Next, in Figure 2c, we study the average size $\langle{s}\rangle$ of shortest paths between connected nodes in the higher-order network as a function of their length $\ell$. We observe that the shortest paths which are longer also tend to be those with a lower average size. This suggests that higher-order interactions are often central in the system, in agreement with recent work on hypercores mancastroppa2024hypercores while dyadic edges remain crucial to connect more peripheral nodes. To further corroborate our intuition, in Figure 2d we plot the fraction of a shortest path that consists of purely dyadic segments, grouped by the length of the path itself. As expected, the curve is increasing, and such a behaviour is again well reproduced in the randomised system.
Figure 3: Summary statistics for shortest paths in static hypergraphs. Average size $\langle{s}\rangle$ as a function of the path length across 25 datasets describing social interactions. (Inset) Average shortest higher-order path length $\langle\ell_{HO}\rangle$ against the average shortest dyadic path length $\langle\ell_{DY}\rangle$ for all datasets.
@@
In Figure 3a we extend the analysis of higher-order shortest paths to all our datasets. We observe that the average size $\langle{s}\rangle$ as a function of $\ell$ decreases in all systems, except for a dataset about social interactions in an office. Finally, in Figure 3b, for all datasets we show the difference in average path length for higher-order and purely dyadic paths. As expected, those systems which have $\langle{s}\rangle\approx 2$ in Figure 3a for all lengths are those which fall closest to the line $x=y$ in Figure 3b. Finally, we observe that the lengths of shortest paths in all datasets are very short, never having a length of more than 5, in agreement with the small-world nature of complex networks wattsstrogatz1998smallworld .
@@
Figure 5: Fastest paths in the temporal hypergraph of the Copenhagen study. Path-duration distributions of higher-order and purely dyadic paths for hypergraphs of the Copenhagen dataset for (a) the empirical data and (b) a randomised null-model. (c) Average interaction size $\langle{s}\rangle$ of the higher-order path as a function of the path duration for empirical (blue) and randomised (dark red) data. (d) Average fraction of each path which is purely dyadic as a function of the path duration for empirical (grey) and randomised (light red) data. Shaded areas represent a 97.5% confidence interval obtained over 100 randomisations.
@@
We model the temporal evolution of a networked system with a temporal network. A temporal network is a sequence of consecutive time-stamped static graphs, or alternatively a collection of nodes and a collection of timestamped edges. At any time $t$ nodes can be active or inactive depending on whether they are participating in an interaction at that instance or not.
@@
A temporal path is a sequence of interactions occurring at successive timestamps. More formally, a temporal path can be defined as an ordered sequence of (node, time) pairs, where consecutive pairs’ timestamps must be increasing. An important consequence of temporality is that paths are not symmetric (even in the case of undirected graphs). In other words, the existence of a path from $i$ to $j$ does not imply the existence of a path from $j$ to $i$ and in general, even when both exist, they might be different and have different lengths. In this work, we assume that a message cannot be propagated further than its direct neighbours within any single snapshot (although the alternative makes sense when timestamps are at longer intervals), and that a node can ‘retain’ a message across multiple snapshots, following the approach of Ref. nicosia2013graphmetrics .
@@
The concept of length in the temporal setting may be defined in 2 complementary ways. The temporal duration, denoted $d$, of a temporal path from node $i$ to node $j$ is defined as the time elapsed between the activation time of node $i$ and the first activation time of node $j$ following the starting time. By contrast, the length or step-count, denoted $\ell$, of a temporal path from nodes $i$ to node $j$ is defined as the total number of segments traversed and is equal to the number of nodes minus one.
@@
Shortest temporal paths can therefore be either those that have minimum duration ($d_{\min}$), or minimum length ($\ell_{\min}$). The fastest path (of shortest duration) between nodes $i$ and $j$ is the temporal path of minimum duration, and is set to infinity if such a path does not exist. The shortest path (with the fewest steps) counts the number of segments, and is hence analogous to the static case.
@@
When a system describes time-varying interactions between groups of nodes, we use temporal hypergraphs. A temporal hypergraph is a sequence of static hypergraphs, indexed by time. The two notions of temporal length (duration and step count) can be defined as in the pairwise case. In Figure 4, a simple temporal hypergraph illustrates how such measures can differ. A temporal path $(A,t_{1})\to(C,t_{1})\to(D,t_{2})\to(D,t_{4})\to(E,t_{4})\to(G,t_{5})$ has a temporal length of 5, but a topological length of 4. In our temporal analysis, we mainly focus on fastest paths, and discuss differences with shortest paths in Figure 7.
@@
We investigate temporal connectivity and fastest paths in various real-world systems to discover how much of the system’s connectivity is due to higher-order interactions. Similarly to the static case, in Figure 5, we begin our exploration of the temporal connectivity properties of real-world systems by focusing on the Copenhagen dataset. In Figure 5a we plot side-by-side the distribution of durations of fastest paths across all interaction sizes in blue, and for only purely pairwise interactions in grey. We note that the duration of such paths can be very long, even well above 100 time units. When only pairwise interactions are considered, there is a distinct distribution shift to the right of temporal path lengths. Such an effect of higher-order interactions is much greater than what we observed for the path length in static hypergraphs, making explicit the very relevant role of non-dyadic connections in determining the connectivity properties of temporal systems.
@@
In Figure 5b we plot the same distribution of durations for a randomisation of the data where hyperedges’ time-information is shuffled but the underlying static graph is kept constant. As shown, such a randomisation fails to reproduce the distribution of durations of the actual system, as washing out temporal correlations among hyperedges makes the typical duration of fastest paths in the null model much shorter and narrower than in the true data.
@@
Next, in Figure 5c we plot the average size $\langle{s}\rangle$ of fastest paths between connected nodes in the higher-order network, grouped by their duration, for both real and randomised data. As in the static case, we observe that paths that last longer also tend to have a lower average size in both the true and randomised networks. Faster paths traverse more higher-order interactions. In contrast, paths with longer duration use more dyadic edges, as illustrated in Figure 5d, which shows the average fraction of fastest paths composed of pure dyadic interactions, grouped by path duration. While the null model qualitatively reproduces the observed trends, we observe that fastest paths in the randomised data systematically underestimate the usage of dyadic edges in temporal paths.
@@
In Figure 6a we broaden our investigation to consider fastest paths for all datasets in our collection and plot $\langle{s}\rangle$ against $t$, the number of timesteps across all datasets. Again, we observe again the same downward trend of decreasing average size with increasing path duration which is even more pronounced when plotting the topological length of fastest paths, and an increasing dyadic proportion for longer path lengths (last two not shown here). In Figure 6b, we compare fastest higher-order path durations $d_{HO}$ with fastest dyadic path durations $d_{DY}$, calculating the percentage of pairs of nodes that are only connected if they may traverse group interactions. While there is variability across real-world systems, we observe that in many cases a high percentage of paths disappear when only true dyadic segments are considered, highlighting the importance of higher-order interactions for the global architecture of the system. For those nodes that remained connected even when only dyadic interactions were allowed, in Figure 6c, we visualise how much longer on average a dyadic temporal path will tend to be than its higher-order counterpart. We note that in a some cases such differences can be very large, such as for the InVS13 and the Baboons datasets.
Datasets $NCC$ $|LCC|$
@@
Finally, in Figure 7 we investigate the differences between path duration between fastest (F) and topologically shortest (S) temporal path, as defined at the beginning of this section. In Figure 7a we plot the percentage of higher-order paths that differ between fastest and shortest for each dataset,. We then determine how much longer shortest paths last than fastest paths by computing the fraction $\langle{d_{\min}^{S}\rangle}/\langle{d_{\min}^{F}}\rangle$ in Figure 7b. As shown, many of the fastest paths are not topologically shortest, with this phenomenon having different intensities in different systems. For instance, in the Copenhagen dataset we reach the highest percentage of different fastest versus topologically shortest temporal paths among pairs of nodes, with this number being close to $60\%$, while in a social hypergraph of interactions among baboons, fastest and shortest paths are almost always the same. In Figure 7b we quantify how much longer shortest paths takes instead of fastest ones across real-world systems by computing the ratio between their average durations. While in some cases topologically shortest paths are on average only a little slower than fastest ones, as in the case of a Sociopatterns data of social interactions in a high school ($d_{\min}^{S}/d_{\min}^{F}=1.1$), in some other cases the duration of shortest paths can be twice as much as that of fastest ones, as for the Congress Bills dataset.
Figure 7: Comparing fastest versus shortest in temporal hypergraphs. For each dataset, we show the percentage of paths that differ between shortest and fastest across all pairs of individuals (a), and how much faster on average the fastest path is when compared to the shortest path (b).

### v3

@@ INTRODUCTION
Networks, collections of nodes and their interactions, are one of the primary tools used to study complex systems, allowing us to describe collective, emergent system properties which can not be captured by looking at the individual units in isolation newman2018networks . A typical case is the emergence of global connectivity from local interactions, with many real-world networks characterised by the emergence of large connected components molloy1995critical . Many such systems display low diameter and are hence dubbed ‘small-worlds’ wattsstrogatz1998smallworld , displaying a relational structure able to support both efficient system-wide communications latora2001efficient ; estrada2008communicability and making them easy to navigate kleinberg2000navigation . The efficient structural connectivity of these systems has a profound impact on their functionality, for both stochastic and deterministic processes. The presence of shortcuts is known to influence the speed of contagion and emergence of cooperation santos2005epidemic , impacting also the ability of a system to synchronise barahona2002synchronization . For all these reasons, efficiently computing shortest paths in graphs is a problem that has attracted enormous interest in the research community for many decades dijkstra2022note , and it continues to do so.
@@
Networks have classically modelled interactions through links, describing relations between pairs of entities only, even though in many real-world systems interactions simultaneously involve multiple nodes battiston2020honetworks . Recently, a wide variety of structural descriptors have been extended to account for the presence of such higher-order interactions, including algorithms for motif discovery lotito2022homotifs and analysis mann2021randomgraphs , community detection eriksson2022flow ; ruggeri2023community , centrality measures benson2019three , as well as network filtering musciotto2021detecting and reconstruction young2021hypergraph procedures. Non-dyadic interactions generate new dynamical behaviours and collective phenomena battiston2021physics ; bick2023higher , from contagion iacopini2019simplicial ; dearruda2020social to synchronisation skardal2020higher ; millan2020explosive ; zhang2023higher and evolutionary games alvarez2021evolutionary ; civilini2024explosive . More recently, temporality has also been considered in higher-order networks, from burstiness cencetti2021temporal and temporal-topological correlations ceria2023temporal , to structural models with gallo2024higher ; iacopini2024temporal and without petri2018simplicial ; di2024percolation memory.
Despite these advances, characterising shortest paths and connectivity in systems with higher-order interactions remains an open problem. Recently, efforts have been devoted to characterise the concepts of distance vasilyeva2023distancesho and walks aksoy2020hypernetwork in networks with non-dyadic ties, as well as proposing efficient algorithms to extract shortest paths in hypergraphs gao2014hoshortestpath , and randomise hypergraphs preserving shortest path lengths najajima2022randomising . However, all these are analyses limited to static systems.
In this work, we provide the first systematic investigation of the effect of higher-order interactions on system connectivity across a variety of real-world datasets of social interactions. We quantify the contribution of non-dyadic ties for shortest path length, computing length distribution, average path size and the fraction of purely dyadic segments in each path, and show that all such features are compatible with a simple null-model which preserves the higher-order degree distribution. Next, we consider time-varying interactions, and extend the concept of higher-order path and components to temporal hypergraphs. Our analysis of multiple real-world systems reveals the crucial role of higher-order interactions to ensure efficient connectivity in temporal higher-order networks, characterises differences between topologically shortest and temporally fastest paths, and shows how the observed non-trivial empirical patterns can not be reproduced with simple randomised null-models. Our work provides new insights into the connectivity properties and the structural organisation of real-world hypergraphs, both for static and time-varying systems.
@@ RESULTS
To study shortest paths in temporal hypergraphs, we collected 24 publicly available datasets of real-world temporal systems with higher-order interactions. Specifically, our study focuses on hypergraphs that describe social interactions from a broad range of domains such as households and schools, hospitals, workspaces and conferences, as well as political interactions between individuals in Congress and even contact data from baboons.
Dataset $V$ $E_{HO}$ $E_{DY}$ $|e_{\max}^{HO}|$ $T$ dtdt
@@
In simple graphs, relational information is described by nodes and edges, which encode pairwise interactions between pairs of nodes only. A path between 2 nodes $i$ and $j$ is an ordered, non-repeating sequence of pairwise edges Similarly, a hypergraph is a collection of nodes and hyperedges which encode interactions among an arbitrary number of nodes. Hyperedges are called pure pairwise or dyadic when they correspond to the classic notion of an edge by connecting only 2 nodes and are distinguished from truly higher-order interactions where 3 or more nodes participate in an interaction. The size of a hyperedge is defined as the number of nodes contained therein. A hyperpath or a higher-order path between 2 nodes in a hypergraph is defined as an ordered, non-repeating sequence of hyperedges, where each subsequent step occurs between nodes within the same hyperedge.
The length of a higher-order path between any two nodes $i$ and $j$, $\ell_{ij}$, is defined as the number of edges traversed. If no such path exists between 2 nodes, we set $\ell_{ij}=\infty$. Two nodes are connected if a path of finite length exists between them. While a pair of nodes can be connected by multiple distinct paths, the shortest path between any two nodes in a static network is defined as a path of minimum length, which might not necessarily be unique. Technical details regarding the exact computation of such hyperpaths are provided in the relevant section of Methods.
Figure 1: A shortest hyperpath between nodes B and K has length $\ell=3$ and average size $\langle{s}\rangle=(4+5+2)/3$.
@@
Clearly, hyperpaths in higher-order networks will traverse hyperedges of varying sizes, where each segment of a hyperpath traverses a hyperedge of a potentially different size. As such, one may attach to each individual segment of a hyperpath its associated size $s$, defined as the size of the hyperedge that is traversed during that particular step. In some cases, such a step may be contained in multiple hyperedges, in which case we choose the hyperedge of smallest size to define its size. Results concerning alternatives, such as selecting the mean or maximum in the case multiple hyperedges, and the frequency with which they occur are reported in Section 3 of the Supplementary Materials. Taken together, our analyses show that different strategies with regard to the hyperedge to traverse in case of redundancies do not induce a qualitative difference in our findings. Paths in higher-order networks contain rich information and one may calculate the average size $\langle{s}\rangle$ of a path as the average size across traversed segments. For example, the average size of a path allows one to determine how much of a path makes use of dyads as opposed to genuinely higher-order ($s\geq 3$) interactions. The minimum size of a path is $\langle{s}\rangle=2$, which occurs when the hyperpath consists of purely pairwise interactions.
@@
We begin by considering static hypergraphs where we neglect information about the temporal nature of the interactions (see Methods for details). As an illustrative example, in Figure 2, we show results for the social hypergraph from the well-known Copenhagen network study, which describes social interactions over 4 weeks between 576 university students on a university campus. Results of analyses on all other datasets are provided in the Supplementary Materials Section 2 and 3.
@@
Next, in Figure 2c, we study the average size $\langle{s}\rangle$ of shortest paths between connected nodes in the higher-order network as a function of their length $\ell$. We observe that shortest paths which are longer also tend to be those with a lower average size. This suggests that higher-order interactions are often central in the system, in agreement with recent work on hypercores mancastroppa2024hypercores while dyadic edges remain crucial to connect more peripheral nodes. Due to the small sample (2) of maximum-length paths, the randomisation does not reflect the downward trend at all lengths. To further corroborate our intuition, in Figure 2d we plot the fraction of a shortest path that consists of purely dyadic segments, grouped by the length of the path itself. As expected, the curve is increasing, and such a behaviour is again well reproduced in the randomised system.
Figure 3: Summary statistics for shortest paths in static hypergraphs. Average size $\langle{s}\rangle$ as a function of the path length across 24 datasets describing social interactions. (Inset) Average shortest higher-order path length $\langle\ell_{HO}\rangle$ against the average shortest dyadic path length $\langle\ell_{DY}\rangle$ for all datasets.
@@
In Figure 3 we extend the analysis of higher-order shortest paths to all our datasets. We observe that the average size $\langle{s}\rangle$ as a function of $\ell$ decreases in all systems, except for a dataset about social interactions in an office. We hypothesize that finite-size effects play a role for this workplace data, as the number of paths of length 3 are three orders of magnitude less than the other path length counts.
The inset of Figure 3 shows for all datasets the difference in average path length for higher-order and purely dyadic paths. As expected, those systems which have $\langle{s}\rangle\approx 2$ in the main figure for all lengths are those which fall closest to the line $x=y$ in the inset. Finally, we observe that the lengths of shortest paths in all datasets are very short, never having a length of more than 5, in agreement with the small-world nature of complex networks wattsstrogatz1998smallworld .
@@
Figure 5: Fastest paths in the temporal hypergraph of the Copenhagen study. Path-duration distributions of higher-order and purely dyadic paths for hypergraphs of the Copenhagen dataset for (a) the empirical data and (b) a randomised null-model. (c) Average temporal interaction size $\langle{s}\rangle$ of the higher-order path as a function of the path duration for empirical (blue) and randomised (dark red) data. (d) Average fraction of each path which is purely dyadic as a function of the path duration for empirical (grey) and randomised (light red) data. Shaded areas represent a 97.5% confidence interval obtained over 100 randomisations.
@@
We model the temporal evolution of a networked system with a temporal network. A temporal network is a sequence of consecutive time-stamped static graphs, or alternatively a collection of nodes and a collection of timestamped edges. At any time $t$ hyperlinks can be either active or inactive.
@@
A temporal path is an ordered sequence of successive (edge, activation time) pairs, where consecutive pairs’ timestamps must be increasing. An important consequence of temporality is that paths are not symmetric (even in the case of undirected graphs). In other words, the existence of a path from $i$ to $j$ does not imply the existence of a path from $j$ to $i$ and in general, even when both exist, they might be different and have different lengths. In this work, we assume that a message cannot be propagated further than its direct neighbours within any single snapshot (although the alternative makes sense when timestamps are at longer intervals), and that a node can ‘retain’ a message across multiple snapshots, following the approach of Ref. nicosia2013graphmetrics .
@@
The concept of length in the temporal setting may be defined in 2 complementary ways. The temporal duration, denoted $d$, of a temporal path from node $i$ to node $j$ is defined as the time elapsed between the timestamp at which node $i$ is first active and the timestamp at which node $j$ is active. By contrast, the length or step-count, denoted $\ell$, of a temporal path from nodes $i$ to node $j$ is defined as the total number of temporal edges traversed.
@@
Shortest temporal paths can therefore be either those that have minimum duration ($d_{\min}$), or minimum length ($\ell_{\min}$). The fastest path (of shortest duration) between nodes $i$ and $j$ is the temporal path of minimum duration, and is set to infinity if such a path does not exist. The shortest path (with the fewest steps) counts the number of steps, and is hence analogous to the static case.
@@
When a system describes time-varying interactions between groups of nodes, we make use of temporal hypergraphs. A temporal hypergraph is a sequence of static hypergraphs, indexed by time. The two notions of temporal length (duration and step count) can be defined as in the pairwise case. In Figure 4, a simple temporal hypergraph illustrates how such measures can differ. A temporal path consisting of timestamped links $(A\to{C},t_{1}),(C\to{D},t_{2}),(D\to{E},t_{4}),(E\to{G},t_{5})$ has a temporal length of 5 units, but a topological length of 4 links. In our temporal analysis, we mainly focus on fastest paths, although we present some results for both fastest and shortest in Section 4 of the Supplementary Materials.
@@
We investigate temporal connectivity and fastest paths in various real-world systems to discover how much of the system’s connectivity is due to temporal higher-order interactions. Similarly to the static case, in Figure 5, we begin our exploration of the temporal connectivity properties of real-world systems by focusing on the Copenhagen dataset. As before, all analysis for the remaining datasets are available in Sections 2 and 3 of the Supplementary Materials, and we mention here only that similar trends are observable across the other datasets. In Figure 5a we plot side-by-side the distribution of durations of fastest paths across all temporal interaction sizes in blue, and for only purely pairwise temporal interactions in grey. We note that the duration of such paths can be very long, even well above 100 time units. When only pairwise interactions are considered, there is a distinct distribution shift to the right of temporal path lengths. Such an effect of higher-order interactions is much greater than what we observed for the path length in static hypergraphs, making explicit the very relevant role of non-dyadic connections in determining the connectivity properties of temporal systems.
@@
In Figure 5b we plot the same distribution of durations for a randomisation of the data where hyperedges’ time-information is shuffled but the underlying static graph is kept constant. As shown, such a randomisation fails to reproduce the distribution of durations of the actual system, as washing out temporal correlations among hyperedges makes the typical duration of fastest paths in the null model much shorter and narrower than in the real-world data.
@@
Next, in Figure 5c we plot the average size $\langle{s}\rangle$ of fastest paths between connected nodes in the higher-order network, grouped by their duration, for both real and randomised data. As in the static case, we observe that paths that last longer also tend to have a lower average size in both the true and randomised networks. Faster paths traverse more higher-order temporal interactions. In contrast, paths with longer duration use more dyadic edges, as illustrated in Figure 5d, which shows the average fraction of fastest paths composed of pure dyadic temporal interactions, grouped by path duration. While the null model qualitatively reproduces the observed trends, we observe that fastest paths in the randomised data systematically underestimate the usage of dyadic edges in temporal paths.
@@
In Figure 6a we broaden our investigation to consider fastest paths for all datasets in our collection and plot $\langle{s}\rangle$ against $t$, the number of timesteps across all datasets. Again, we observe again the same downward trend of decreasing average size with increasing path duration which is even more pronounced when plotting the topological length of fastest paths, and an increasing dyadic proportion for longer path lengths (last two not shown here). In Figure 6b, we calculate the proportion of node-pairs in each dataset that remain connected when dyadic edges are removed, i.e. when they are only allowed to traverse temporal interactions involving groups. While there is variability across real-world systems, we observe that in many cases a high percentage of paths disappear when only pure dyadic segments are considered, highlighting the importance of higher-order interactions for the global architecture of the system. For those nodes that remained connected even when only dyadic temporal interactions were allowed, in Figure 6c, we visualise how much longer on average a dyadic temporal path will tend to be than its higher-order counterpart. We note that in a some cases such differences can be very large, such as for the InVS13 and the Baboons datasets.
Datasets NCCNCC $|LCC|$